the slope of a line perpendicular to the line containing (-6,1) and (3,-2)

Answer by jpvn2015(54)

First find the slope of the two points: y2-y1 / x2-x1 = slope
-2-1 / 3--6 = -3 / 3+6 = -3/9 = -1/3
Thus, the slope of the original line is -1/3.
To find the perpendicular slope, all you have to do is take the reciprocal of -1/3, which is -3, and multiply it by "-1."
-3 x -1 = 3. Thus, your answer would be 3.